OpenStudy (anonymous):

Can anyone help out this question....? Arrange the following functions in increasing order of growth rate. ie. if g(n) follows f(n) in your list, then f(n) is necessarily O(g(n)). a)2^2n b)2^n^2 c)n^2log(n) d)n e)n^2n

OpenStudy (anonymous):

OK, what do you think?

OpenStudy (anonymous):

you have \[2^{2*n}\] and \[2^{n^2}\] and \[n^{2 * log(n)}\] and \[n\] and \[n^{2*n}\]

OpenStudy (anonymous):

For this you must imagine \[n\] increasing

OpenStudy (anonymous):

You should graph this first

OpenStudy (anonymous):

that is an easier way to see how the functions change as \[n\] increases

OpenStudy (anonymous):

third option is \[n^{2}\log(n)\]

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

well, which do you think is smallest, therefore the one that does not change fastest

OpenStudy (anonymous):

n will be

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

so compare \[2^{2*n}\] and \[2^{n^2}\]

OpenStudy (anonymous):

because they have the same base, we can more easily compare them. Which would be greater than the other as n increases?

OpenStudy (anonymous):

2^2n will be smaller

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Now compare \[n^{2}*log(n)\] and \[n^2*n\]

OpenStudy (anonymous):

which is larger

OpenStudy (anonymous):

as n increases

OpenStudy (anonymous):

its n^2^n

OpenStudy (anonymous):

a)2^2n b)2^n^2 c)n^2log(n) d)n e)n^2n

OpenStudy (anonymous):

\[2^{2*n}\] \[2^{n^2}\] \[n^{2} * log(n)\] \[n\] \[n^{2*}\]

OpenStudy (anonymous):

\[n^{2*n\]

OpenStudy (anonymous):

option e is n^2^n

OpenStudy (anonymous):

\[n^{2^{n}}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[2^{2*n}\] \[2^{n^{2}}\] \[n^2*log(n)\] \[n\] \[n^{2^{n}}\]

OpenStudy (anonymous):

so think n^2^n will be larger than n^ log(n)

OpenStudy (anonymous):

so know we had arrange in increasing of their growth rate ie. if g(n) follows f(n) , then f(n) is necessarily O(g(n)).

OpenStudy (anonymous):

OK, yes, big-o notation

OpenStudy (anonymous):

so what will be answer?

OpenStudy (anonymous):

so you have some relative ordering of functions, now you can compare them by equating them, because then, any number beyond that point of equation will help you determine which function is greater

OpenStudy (anonymous):

Something are best done with pen and paper so you can write out your work then upload an image you want

OpenStudy (anonymous):

n< 2^2n < 2^n^2 < n^2 log(n) < n^2n

OpenStudy (anonymous):

is it correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hey, I didn't check it yet, was AFK

OpenStudy (anonymous):

OK, so

OpenStudy (anonymous):

we know n is the smallest

OpenStudy (anonymous):

However, equate \[2^{2*n}\] and \[n^{2} *log(n)\]

OpenStudy (anonymous):

n^2 gets beaten by 2^n

OpenStudy (anonymous):

what is 100^2 and what is 2^100? the first is just 100 * 100, then next is 2 * 2 * 2 * ... * 2 100 times

OpenStudy (anonymous):

110^2 = 10000 and 2^100 = 1.2676506 × (10^30) which is much bigger than 10,000

OpenStudy (anonymous):

understood

OpenStudy (anonymous):

so, you just need to plug in very large numbers and compare all equations

OpenStudy (anonymous):

n^2 log(n) will be larger

OpenStudy (anonymous):

essentially

OpenStudy (anonymous):

yes, but only by the factor of log(n)

OpenStudy (anonymous):

so 100^2 * log(100) is just 20,000

OpenStudy (anonymous):

log is slow moving function

OpenStudy (anonymous):

value log will be always positive so it increases

OpenStudy (anonymous):

by factor of log(n)

OpenStudy (anonymous):

when does n^2 * log(n) = 2^(2*n) ?

OpenStudy (anonymous):

when n=0

OpenStudy (anonymous):

OK, cool, so when n = 1, which is larger? I choose 1 because in order to see which function is larger we must increase n now that they are equal

OpenStudy (anonymous):

continue this method of comparison

OpenStudy (anonymous):

for all functions